OpenStudy (anonymous):

1+3+5+...+499 find the sum

OpenStudy (anonymous):

If its just adding then its 508.

OpenStudy (marsxtc):

sum = n(n+1)/2 n = highest number there.

OpenStudy (marsxtc):

Tell me if you need more help.

OpenStudy (anonymous):

it says use S=n^2 to find the sum but ive tried it a million times

OpenStudy (marsxtc):

I'm not sure then. I would just use the equation I gave you.

OpenStudy (anonymous):

I did and it's not any of the answers on the review sheet I got.